find a general equation for energy difference between two consecutive
state for a particle confined within a potential width of a
1
Expert's answer
2019-09-16T09:47:49-0400

Since energy is at the level n "E_n = \frac {h^{2}n^{2}}{8ml^{2}}"


then the energy difference "\Delta E_n = E_{n+1} - E_n = \frac{h^{2}(2n+1)}{8ml^{2}}"